# Pulsed Vector Atomic Magnetometer Using an Alternating Fast-Rotating Field

### v3

@@ II.3.1 Results of the systematics
$|B_{PD}|$ ($\theta\approx 21^{\circ}$, $\beta\approx 0$) 2.6 nT 2.8 nT
[Berry’s phase] The Berry phase will result in a phase shift that adds to the total magnetic field measurement. When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq. 6 predicts 4.7 nT ($\phi_{B}/\gamma$), as shown in Table. 2.
@@
[Second harmonic phase shift] The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 7 prediction well.
@@
[Heading Errors] The static component of the dynamic heading error is measured when the magnetometer sensor rotates by an angle $\beta=24^{\circ}$ in the yz-plane, resulting in a heading error of $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$. From Eq. (8), the static heading error is estimated to be 2.9 nT, with an experimental result of 2.7 nT. The slightly lower experimental value is attributed to incomplete spin polarization.
(a) Probe beam heading error
(b) Dynamic heading error
(c) Eddy current due to different switches
(d) Eddy current time constant
@@
Figure 2: (a) The measurement of the probe beam heading error. The solid points are experimental results, and the dashed lines are the theoretic plots based on our proposed model. (b) The magnetic field in the transverse direction is caused by the eddy current. The magnetic fields measured by both the Zero-Altering and Peak-Altering are subtracted by 16.6 nT to remove the residual magnetic field.(c) The measurement of the eddy current induced magnetic field due to different switches. (d) The measurement of the eddy current time constant.
@@
A rotating magnetic field can induce a rotating-field-phase-dependent heading error, $B_{PD}$. This effect is characterized in Table 3 with both experimental results and density matrix simulations [32]. The density matrix model, which represents an ensemble of spins in a mixed state, is described in detail in Secs. VIII and IX of the supplemental material.
When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be interpreted as the spin precession plane adiabatically following the rotation of the total magnetic field $\mathbf{B_{tot}}$. This keeps the relative angle between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a steady precession frequency. Consequently, the dynamic heading error depends on the initial angle between the spin polarization and $\mathbf{B_{tot}}$, which remains equal to its initial value. We define $\Theta$ as the angle by which $\mathbf{B_{tot}}$ deviates from the position where $\mathbf{S}$ is perpendicular to $\mathbf{B_{tot}}$, as shown in Fig. 2(b).
For example, when $\mathbf{B_{m}}$ begins along the x-axis, the spins are polarized in a direction tilted by an angle $\beta$ away from the negative y-axis in the yz-plane, and $\mathbf{B_{tot}}$ initially lies in the xz-plane. Due to symmetry, whether $\mathbf{B_{m}}$ starts from the positive or negative x-axis, the relative angles between $\mathbf{S}$ and $\mathbf{B_{tot}}$ remain the same. In this configuration, the dynamic heading error equals the static heading error, $\mathcal{B_{H}}\cos\theta\sin\beta$, with $B_{PD}=0$. Under these conditions, the spin precession plane follows the rotation of the total magnetic field while keeping $\Theta$ constant. Specifically, if $\beta=0$ initially, $\mathbf{S}$ stays perpendicular to $\mathbf{B_{tot}}$ throughout, resulting in $B_{DH}=0$.
However, if $\mathbf{B_{m}}$ originates from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the yz-plane, then $\mathbf{S}$ deviates from perpendicularity with $\mathbf{B_{tot}}$ by an angle $\Theta=\beta+\theta$. This results in a dynamic heading error of $\mathcal{B_{H}}\sin(\beta+\theta)$, with $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta\approx 2.9$ nT, and $B_{DH}=\mathcal{B_{H}}\cos\beta\sin\theta\approx 2.5~{}\mathrm{nT}$. Conversely, if the rotating field begins from the positive y-axis, $\Theta=\beta-\theta$, with $B_{SH}$ remaining the same, but $B_{DH}=-\mathcal{B_{H}}\cos\beta\sin\theta\approx-2.5~{}\mathrm{nT}$. Table 1 provides a full overview of the sign dependence of the dynamic heading error.
@@
Table 3: Measured dynamic heading error along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$. Here the static part of dynamic heading error $B_{SH}=0$. “Exp” is the experimental data, “DM” is the density matrix simulation result.
$\mathbf{B_{m}}$ initiates from Y-axis $\mathbf{B_{m}}$ initiates from X-axis ($\mathbf{B_{tot}}\perp\mathbf{S}$)
$B_{tot}$ Shot No. Exp. (nT) DM (nT) Exp. (nT) DM (nT)
@@
The experimental results for the dynamic heading error are smaller than the analytical model and density matrix simulation because the actual spin polarization is smaller than the simulation, which assumes full polarization. For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Sec. IX of the supplemental material.
@@
The variation in $B_{x}$ measurement from the density matrix simulation in each shot is due to the probe beam heading error ($\phi_{\text{prob}}/\gamma$). For example, if the sensor is misaligned with an angle of $\alpha=1^{\circ}$, this results in an approximate 1.2 nT probe beam heading error in the x-direction. Additional data in Sec. IX of the supplemental material indicate that these magnetic field offsets maintain their signs even when the polarization direction is reversed, confirming that they are not caused by dynamic heading error. These systematic errors, however, can be eliminated through 4-shot averaging.
@@
To investigate the probe beam heading error, the sensor rotates about the y-axis within the xz-plane at an angle of $\alpha$. This error can introduce a systematic effect in transverse magnetic field measurements that depends on the angle $\alpha$, with its sign determined by the rotation direction of the rotating field. The measured probe heading error at $\omega_{m}=2\pi\times 480$ Hz is shown in Fig.2(a), with the theoretical basis provided in Eq.9. “Theory 1” corresponds to anti-clockwise rotations (Shots 1 and 3 in Table 1), given by $-\omega_{m}/\gamma\cdot\tan\alpha$, while “Theory 2” represents clockwise rotations (Shots 2 and 4 in Table 1), given by $\omega_{m}/\gamma\cdot\tan\alpha$. The probe beam heading error can be effectively canceled by averaging the measurements from two opposite rotating field directions.
@@
[Eddy current] The alternating fast rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield. In our four-shot scheme, which alternates the rotation direction, there are two switches within four shots, as shown in Fig.1b. We found that the way the rotating magnetic field is altered affects the eddy current and its induced magnetic field. When the direction of the rotating field changes at the peak values of the field—from negative maximum to positive maximum and vice versa—we refer to this as “Peak-Altering” (Fig.1b, top left). In this case, the average magnetic field caused by the eddy current equals zero. However, if the direction change occurs when the rotating magnetic field is at zero, termed “Zero-Altering” (Fig. 1b, top right), it leads to a significant magnetic field caused by the eddy current.
@@
In our setup, the y-direction switch of $B_{y}$ occurs during the preparation times of the second and fourth shots, as shown in Fig.1b. In the “Zero-Altering” case, the second and fourth shots experience the maximum eddy current, while the first and third shots have smaller eddy currents, as they capture the decaying eddy current generated during the fourth and second shots. The sign of the eddy current-induced magnetic field depends on the rotation direction of the rotating field. A panorama measurement (80 shots) result from the MDA, shown in Fig.2(c), indicates that Zero-Altering can cause an eddy current magnetic field of approximately 450 nT. In contrast, Peak-Altering effectively suppresses the systematic effect of the eddy current. Any remaining small eddy current magnetic field in Peak-Altering is due to imperfections, as instruments have a finite rise time during abrupt signal transitions.
@@
To characterize the time constant of the eddy current, we applied a continuous rotating magnetic field with a constant phase (no alteration) over one panorama measurement. In Fig. 2(d), 79 data points represent 79 individual shot measurements, with the first shot discarded due to phase shift overload in the MDA. Each shot measurement lasts 1/120 s. The 60 Hz line frequency produces periodic variations and harmonics, which are effectively averaged out over one panorama. Our setup’s measured eddy current time constant is approximately 10.4 ms, corresponding to a cut-off frequency of 15 Hz.
@@
To mitigate the eddy current’s systematic effect, we designed a waveform shown in Fig. 1b to ensure that rotation direction alterations consistently occur at the peak values of the rotating magnetic field (Peak-Altering). By employing Peak-Altering and 4-shot averaging, we effectively reduce the eddy current’s systematic effect.
II.3.2 Results of magnetic field sensitivities
@@
A magnetic field gradiometer measures the variation in magnetic field strength over a short distance, rather than the absolute field strength at a single location. This method is effective for detecting local magnetic changes while reducing the influence of uniform, distant magnetic interference [39].To demonstrate the gradiometric measurement capability of the FRF vector magnetometer, we configured a setup using a multipass 87Rb cell and a high-power QPC laser with a 25 W output. The laser, operating in pulsed mode, polarizes the spins along the z-axis to enhance spin alignment and is then turned off during the readout phase. Following this pumping stage, $\pi/2$ magnetic field pulses are applied to tip the spins from the z-axis to the y-axis.
@@
Figure 3: Magnetic gradiometer sensitivity: The solid lines represent the sensitivities of the FRF vector magnetometer for the total magnetic field and the two transverse magnetic fields. The dashed line indicates the sensitivity of the magnetometer when operating as a scalar magnetometer. The inset shows the frequency response for $B_{x}$ and $B_{y}$.
III Discussion
@@
IV Methods
IV.1 Shot noise limit
To derive the shot noise limit for the FRF magnetometer, we utilize the CRLB [41] along with photon shot noise considerations specific to balanced polarimeter detection [42]. Based on these principles, the power spectral density $\rho(\omega)$ achieves its minimum when the measurement duration $t\approx 2T_{2}$ (further details in Sec. IV of the supplemental material). For total magnetic field measurements, the sensitivity is expressed as
@@ IV.1 Shot noise limit
If the measurement time matches the optimal duration for total magnetic field measurement ($t=2T_{2}$), the sensitivity for transverse magnetic field measurements is
@@
These derived sensitivities assume independent fitting of scalar and transverse magnetic fields. However, this assumption becomes inaccurate if the modulation frequency is not sufficiently high. Analytical sensitivity solutions can be obtained by simultaneously fitting scalar and transverse magnetic fields using the Minimum-Variance Unbiased (MVU) estimator, as detailed in Sections III and IV of the supplemental material. The results show that when modulation frequency is sufficiently fast ($\omega_{m}\gtrsim\pi/T_{2}$), independent fitting yields comparable sensitivities.
@@
To compare, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers is given by (refer to Sec. VI of the supplemental material for details):
@@ Acknowledgements
The authors thank Dmitry Budker, Conlon Lorcan Oneill for the valuable comments. The authors acknowledge support from the Defense Advanced Research Projects Agency (DARPA), A*STAR Career Development Fund ( [^2] D800028), Italy-Singapore science and technology collaboration grant (R23I0IR042), Delta-Q (C230917004, Quantum Sensing), and Competitive Research Programme (NRF-CRP30-2023-0002).
Author Contributions
T.W. prepared the setup and conducted experiments for the FRF vector magnetometers, proposed and implemented the modulation, modeled spin evolution using rotation and density matrices, derived fundamental limits, analyzed results, and prepared the manuscript. M.R. supervised the setup and experiments, contributed to spin evolution modeling with rotation matrices, and discussed modulation and dynamic heading error. W.L. assisted with modeling static heading error and analyzing atomic polarization. M.L., T.K., and E.F. provided scalar magnetometer sensors, offered valuable experimental advice, and provided constructive comments for manuscript revisions. All authors reviewed the manuscript.
Competing interests
The authors declare no competing interests.

### v4

@@ 2.3.1 Results of the systematics
$|B_{P\!D}|$ ($\theta\approx 21^{\circ}$, $\beta\approx 0$, $B_{m}$ initiates from Y) 2.6 nT 2.8 nT
@@
The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 7 prediction well.
@@
The static component of the dynamic heading error is measured when the magnetometer sensor rotates by an angle $\beta=24^{\circ}$ in the yz-plane, resulting in a heading error of $B_{S\!H}=\mathcal{B_{H}}\cos\theta\sin\beta$. From Eq. (8), the static heading error is estimated to be 2.9 nT, with an experimental result of 2.7 nT. The slightly lower experimental value is attributed to incomplete spin polarization. The predicted value for $|B_{P\!D}|$ is based on approximate analytical model detailed in Supplementary Table 3
@@
Figure 4: Plots of systematic measurements: (a) Dynamic heading error: The spin precession plane adiabatically follows $\mathbf{B_{tot}}$. (b) Measurement of the probe beam heading error: The solid points represent experimental results, while the dashed lines correspond to theoretical predictions based on the proposed model. (c) Eddy currents generated by different rotating field alterations: Measurement of the transverse magnetic field (y-axis) induced by eddy currents due to zero-altering and peak-altering modulations. (d)Eddy Current Time Constant: Measurement of the eddy current time constant during a single panorama measurement of MDA.
@@
A rotating magnetic field can induce a rotating-field-phase-dependent heading error, $B_{P\!D}$. This effect is characterized in Table 3 with both experimental results and density matrix simulations [32]. The density matrix model, which represents an ensemble of spins in a mixed state, is described in detail in Supplementary Note 8 and Supplementary Tables of the supplemental Information.
@@
Table 3: Measured dynamic heading error, primarily influenced by $B_{PD}$, along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$. Here the static part of dynamic heading error $B_{SH}\approx 0$. “Exp” is the experimental data, “DM” is the density matrix simulation result.
$\mathbf{B_{m}}$ initiates from Y $\mathbf{B_{m}}$ initiates from X ($\mathbf{B_{tot}}\perp\mathbf{S}$)
$B_{tot}$
Shot
No.
Exp
DM
Exp
DM
@@
When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be interpreted as the spin precession plane adiabatically following the rotation of the total magnetic field $\mathbf{B_{tot}}$. This keeps the relative angle between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a steady precession frequency. Consequently, the dynamic heading error depends on the initial angle between the spin polarization and $\mathbf{B_{tot}}$, which remains equal to its initial value. We define $\Theta$ as the angle by which $\mathbf{B_{tot}}$ deviates from the position where $\mathbf{S}$ is perpendicular to $\mathbf{B_{tot}}$, as shown in Fig. 4a.
For example, when $\mathbf{B_{m}}$ begins along the x-axis ($\phi_{x}=\pi/2$, $\phi_{y}=0$), the spins are polarized in a direction tilted by an angle $\beta$ away from the negative y-axis in the yz-plane, and $\mathbf{B_{tot}}$ initially lies in the xz-plane. Due to symmetry, whether $\mathbf{B_{m}}$ starts from the positive or negative x-axis, the relative angles between $\mathbf{S}$ and $\mathbf{B_{tot}}$ remain the same. In this configuration, the dynamic heading error equals the static heading error, $\mathcal{B_{H}}\cos\theta\sin\beta$, with $B_{P\!D}=0$. Under these conditions, the spin precession plane follows the rotation of the total magnetic field while keeping $\Theta$ constant. Specifically, if $\beta=0$ initially, $\mathbf{S}$ stays perpendicular to $\mathbf{B_{tot}}$ throughout, resulting in $B_{D\!H}=0$.
However, if $\mathbf{B_{m}}$ originates from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the yz-plane, then $\mathbf{S}$ deviates from perpendicularity with $\mathbf{B_{tot}}$ by an angle $\Theta=\beta+\theta$. This results in a dynamic heading error of $\mathcal{B_{H}}\sin(\beta+\theta)$, with $B_{S\!H}=\mathcal{B_{H}}\cos\theta\sin\beta$, and $B_{P\!D}=\mathcal{B_{H}}\cos\beta\sin\theta$. Conversely, if the rotating field begins from the positive y-axis, $\Theta=\beta-\theta$, with $B_{S\!H}$ remaining the same, but $B_{P\!D}=-\mathcal{B_{H}}\cos\beta\sin\theta$. Table 1 provides a full overview of the sign dependence of the dynamic heading error.
The experimental results for the dynamic heading error are smaller than the analytical model and density matrix simulation because the actual spin polarization is smaller than the simulation, which assumes full polarization. For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Supplementary Tables of the Supplemental Information.
@@
The variation in $B_{x}$ measurement from the density matrix simulation in each shot is due to the probe beam heading error ($\phi_{\text{prob}}/\gamma$). For example, if the sensor is misaligned with an angle of $\alpha=1^{\circ}$, this results in an approximate 1.2 nT probe beam heading error in the x-direction. Additional data in Supplementary Tables of the Supplemental Information indicate that these magnetic field offsets maintain their signs even when the polarization direction is reversed, confirming that they are not caused by dynamic heading error. These systematic errors, however, can be eliminated through 4-shot averaging.
@@
To investigate the probe beam heading error, the sensor rotates about the y-axis within the xz-plane at an angle of $\alpha$. This error can introduce a systematic effect in transverse magnetic field measurements that depends on the angle $\alpha$, with its sign determined by the rotation direction of the rotating field. The measured probe heading error at $\omega_{m}=2\pi\times 480$ Hz is shown in Fig. 4b, with the theoretical basis provided in Eq. 9. “Theory 1” corresponds to anti-clockwise rotations (Shots 1 and 3 in Table 1), given by $-\omega_{m}/\gamma\cdot\tan\alpha$, while “Theory 2” represents clockwise rotations (Shots 2 and 4 in Table 1), given by $\omega_{m}/\gamma\cdot\tan\alpha$. The probe beam heading error can be effectively canceled by averaging the measurements from two opposite rotating field directions.
@@
The alternating fast rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield. In our four-shot scheme, which alternates the rotation direction, there are two switches within four shots, as shown in Fig. 2. We found that the way the rotating magnetic field is altered affects the eddy current and its induced magnetic field. When the direction of the rotating field changes at the peak values of the field—from negative maximum to positive maximum and vice versa—we refer to this as “Peak-Altering” (Fig. 2, top left). In this case, the average magnetic field caused by the eddy current equals zero. However, if the direction change occurs when the rotating magnetic field is at zero, termed “Zero-Altering” (Fig. 2, top right), it leads to a significant magnetic field caused by the eddy current.
@@
In our setup, the y-direction switch of $B_{y}$ occurs during the preparation times of the second and fourth shots, as shown in Fig. 2. In the “Zero-Altering” case, the second and fourth shots experience the maximum eddy current, while the first and third shots have smaller eddy currents, as they capture the decaying eddy current generated during the fourth and second shots. The sign of the eddy current-induced magnetic field depends on the rotation direction of the rotating field. A panorama measurement (80 shots) result from the MDA, shown in Fig. 4c, indicates that Zero-Altering can cause an eddy current magnetic field of approximately 450 nT. In contrast, Peak-Altering effectively suppresses the systematic effect of the eddy current. Any remaining small eddy current magnetic field in Peak-Altering is due to imperfections, as instruments have a finite rise time during abrupt signal transitions.
@@
To characterize the time constant of the eddy current, we applied a continuous rotating magnetic field with a constant phase (no alteration) over one panorama measurement. In Fig. 4d, 79 data points represent 79 individual shot measurements, with the first shot discarded due to phase shift overload in the MDA. Each shot measurement lasts 1/120 s. The 60 Hz line frequency produces periodic variations and harmonics, which are effectively averaged out over one panorama. Our setup’s measured eddy current time constant is approximately 10.4 ms, corresponding to a cut-off frequency of 15 Hz.
@@
To mitigate the eddy current’s systematic effect, we designed a waveform shown in Fig. 2 to ensure that rotation direction alterations consistently occur at the peak values of the rotating magnetic field (Peak-Altering). By employing Peak-Altering and 4-shot averaging, we effectively reduce the eddy current’s systematic effect.
2.3.2 Results of magnetic field sensitivities
@@
A magnetic field gradiometer measures variations in magnetic field strength over a short distance, rather than the absolute field strength at a single location. This approach is highly effective for detecting local magnetic changes while minimizing the influence of uniform, distant magnetic interference [39]. To demonstrate the gradiometric measurement capability of the FRF vector magnetometer, we configured a setup using a multipass 87Rb cell and a high-power QPC laser with a 25 W output. The laser operates in pulsed mode to polarize the spins along the z-axis, enhancing spin alignment, and is turned off during the readout phase. After the pumping stage, $\pi/2$ magnetic field pulses are applied to tip the spins from the z-axis to the y-axis. Notably, for synchronized optical pumping that polarizes spins approximately perpendicular to the external magnetic field, the average power requirement is typically only in the tens of milliwatts, while achieving comparable performance.
@@
Figure 5: Magnetic gradiometer and polar angle sensitivities: The solid lines illustrate the sensitivities of the FRF vector magnetometer for the total magnetic field and the two transverse magnetic fields. The dashed line depicts the sensitivity of the magnetometer in scalar mode. The inset highlights the frequency response for $b_{x}$ and $b_{y}$. The polar angle sensitivities, applicable only to $b_{x}$ and $b_{y}$, are shown on the right-side y-axis and are calculated by dividing $b_{x}$ and $b_{y}$ by the total static magnetic field, $B_{0}$.
3 Discussion
@@
4 Methods
4.1 Shot noise limit
To derive the shot noise limit for the FRF magnetometer, we utilize the CRLB [41] along with photon shot noise considerations specific to balanced polarimeter detection [42]. Based on these principles, the power spectral density $\rho(\omega)$ achieves its minimum when the measurement duration $t\approx 2T_{2}$ (further details in Supplementary Note 4 of the Supplemental Information). For total magnetic field measurements, the sensitivity is expressed as
@@ 4.1 Shot noise limit
When the measurement time matches the optimal duration for total magnetic field measurement ($t=2T_{2}$), the sensitivity for transverse magnetic field measurements is
@@
These derived sensitivities assume independent fitting of scalar and transverse magnetic fields. However, this assumption becomes inaccurate if the modulation frequency is not sufficiently high. Analytical sensitivity solutions can be obtained by simultaneously fitting scalar and transverse magnetic fields using the Minimum-Variance Unbiased (MVU) estimator, as detailed in Supplementary Notes 4 and 5 of the Supplemental Information. The results show that when modulation frequency is sufficiently fast ($\omega_{m}\gtrsim\pi/T_{2}$), independent fitting yields comparable sensitivities.
@@
To compare, the fundamental sensitivity of the transverse magnetic field for sequential modulation (SM) vector magnetometers is given by (refer to Supplementary Note 6 of the Supplementary Information for details):
@@
Data availability
All data necessary to evaluate the conclusions of this study are included in the article, with supplementary data provided in tables and additional information included in the Supplementary Information. Any further details related to this study can be requested from the corresponding authors.
@@ Acknowledgements
The authors thank Dmitry Budker and Conlon Lorcan Oneill for their valuable comments. All authors acknowledge support from the Defense Advanced Research Projects Agency (DARPA) under contract No. HR001120C009 (M.R., T.K.). T.W. acknowledges support from A*STAR Career Development Fund ( [^2] D800028)(T.W.), Italy-Singapore science and technology collaboration grant (R23I0IR042)(T.W.), Delta-Q (C230917004, Quantum Sensing)(T.W.), and Competitive Research Programme (NRF-CRP30-2023-0002)(T.W.).